OpenStudy (anonymous):

give a equation whose solution set is the line joining (2,3) and (-4,6). what is its equation in slope-intercept form what are its x and y intercepts

OpenStudy (anonymous):

First we have to get the slope using the slope formula for "m"\[m = \frac{ y _{1} - y _{2} }{ x _{1} - x _{2} }\]So, m = (6 - 3)/(-4 - 2) = -1/2 Next, we use the point-slope formula:\[y - y _{1} = m(x - x _{1})\]y - 6 = (-1/2)(x + 4) And lastly, we put this into slope-intercept form: y = (-1/2)x + 4 The intercepts are: (0, 4) and (8, 0)

OpenStudy (anonymous):

All good now @ttford8142 ?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @ttford8142